OpenStudy (anonymous):

A train begins a trip of 240 miles. The train averages 40 miles per hour including stops. Create a table, a graph, and an equation to represent the distance the train travels in relation to time.

OpenStudy (anonymous):

remains distance = 240 - 40(hours passed)

OpenStudy (anonymous):

distance=240 times hours passed

OpenStudy (anonymous):

s= vt

OpenStudy (anonymous):

Itss in d format I have 2 options: d=t/40 d= 40t = have chosen this one but not sure

OpenStudy (anonymous):

D=40t

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

s=vt v is velocity that is 40 miles per hour

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

second one is correct

OpenStudy (anonymous):

Thank you Roya

OpenStudy (anonymous):

it gives you passed distance

OpenStudy (anonymous):

do you know whY ?

OpenStudy (anonymous):

yes she knows